OpenStudy (anonymous):

R(p)=1200p=25p^2. Find the unit price that will yield maximum revenue. What is the maximum revenue?

OpenStudy (gorv):

i think it should be R(p)=1200+25^2 ???

OpenStudy (gorv):

helloooo @yoitsren waiting for u to rplyyyy buddy

OpenStudy (anonymous):

sorry, I was away for awhile! I wrote the equation the way it was written in my assignment

OpenStudy (gorv):

r(p)=1200p it means it can generate infinite amount of revenue

OpenStudy (gorv):

1200p=25p^2 1200p/25=p^2 48p=p^2 p^2-48p=0 differentiate it 2p-48=0 p=48/2=24 p=24

OpenStudy (gorv):

at this value revenue will be max